OpenStudy (anonymous):

What is y=2x^2-4x+12 in vertex form?

OpenStudy (jack1):

y= a(x-h)^2+k from y = ax^2 + bx + c

OpenStudy (anonymous):

Help me on my question please

OpenStudy (jack1):

so complete the square of your equation y=2x^2-4x+12

OpenStudy (dan815):

do complete tjhe square

OpenStudy (dan815):

deww it

OpenStudy (dan815):

harry potter is helping u

OpenStudy (jack1):

@Mitch2014 ... u heard the dan

OpenStudy (jack1):

hahaha, someone finally noticed the scar! nice!